Monica and Max Gordon each want to buy a scooter. Monica has already saved twenty-five dollars and plans to save five dollars per week until she can buy the scooter. Max has sixteen dollars and plans to save eight dollars per week.
1. In how many weeks will Monica and Max have saved the same amount of money?
2. How much will each person have saved at that time?

Answer by rfer(16322)

25+5x=16+8x
9=3x
x=3 wks
25+5*3=16+8(3)
40=40
$40